# Fragment-Masked Molecular Optimization

### v2

@@ Abstract
Molecular optimization is a crucial aspect of drug discovery, aimed at refining molecular structures to enhance drug efficacy and minimize side effects, ultimately accelerating the overall drug development process. Many target-based molecular optimization methods have been proposed, significantly advancing drug discovery. These methods primarily on understanding the specific drug target structures or their hypothesized roles in combating diseases. However, challenges such as a limited number of available targets and a difficulty capturing clear structures hinder innovative drug development. In contrast, phenotypic drug discovery (PDD) does not depend on clear target structures and can identify hits with novel and unbiased polypharmacology signatures. As a result, PDD-based molecular optimization can reduce potential safety risks while optimizing phenotypic activity, thereby increasing the likelihood of clinical success. Therefore, we propose a fragment-masked molecular optimization method based on PDD (FMOP). FMOP employs a regression-free diffusion model to conditionally optimize the molecular masked regions without training, effectively generating new molecules with similar scaffolds. On the large-scale drug response dataset GDSCv2, we optimize the potential molecules across all 945 cell lines. The overall experiments demonstrate that the in-silico optimization success rate reaches 94.4%, with an average efficacy increase of 5.3%. Additionally, we conduct extensive ablation and visualization experiments, confirming that FMOP is an effective and robust molecular optimization method. The code is available at: https://anonymous.4open.science/r/FMOP-98C2.
Introduction
Molecular optimization plays a crucial role in drug discovery (Xia et al. 2024), which involves the modification and improvement of lead compounds identified through initial screening to enhance their drug-like properties (Wu et al. 2022). Historically, molecular optimization is planned manually according to knowledge and experience in the pharmacology, and optimized through fragment-based screening or synthesis (Chen et al. 2021). However, manual molecular optimization is not easily scalable to different needs and cannot be automated for large-scale optimization; thus, this strategy is insufficient for meeting the demands of current drug discovery (Schneider et al. 2020).
In recent years, deep learning (DL) methods, particularly diffusion models (Ho, Jain, and Abbeel 2020), have been observed to effectively optimize molecules that meet specific conditions (Zhou et al. 2024; Gu et al. 2024), with the potential to accelerate traditional paradigms. Molecular optimization can rapidly identify potential drug candidates using existing experimental data and molecular structures, reducing the need for blind experimentation and enhancing research efficiency (Choudhuri et al. 2023). Several molecular optimization methods have been proposed to enhance candidate molecule performance, such as target pocket- (Huang et al. 2024) and property-based molecular optimization (Lee et al. 2024).
@@ Introduction
Target pocket-based molecular optimization methods rely on understanding specific pocket structures and their hypothesized roles in combating diseases (Moffat et al. 2017). Challenges such as a limited number of available targets and difficulty capturing clear structures hinder innovative drug development. In contrast, phenotypic drug discovery (PDD) (Vincent et al. 2022a) does not depend on well-defined target structures and can identify hit compounds with novel and multi-target properties. PDD emphasizes the phenotypic effects of molecules within disease-related biological systems (Sadybekov and Katritch 2023) and has significantly contributed to the discovery of first-in-class drugs (Swinney 2013). By observing the phenotypic changes that molecules induce in cells, tissues, or organisms, PDD identifies potential drugs without requiring prior knowledge about specific targets (Vincent et al. 2022b). As a result, PDD’s potential as a drug discovery tool that addressing complex diseases that are not yet fully understood by the scientific community is evident. Extensive PDD research has been conducted, with significant efforts made in constructing relevant datasets such as the genomics of drug sensitivity in cancer (GDSC) (Yang et al. 2013; Garnett et al. 2012). Based on these datasets, numerous artificial intelligence-driven methods (Hostallero et al. 2023; Li et al. 2024b, c) have been proposed to predict drug responses against specific cell lines, typically quantified by the half-maximal inhibitory concentration (IC50) (Cho et al. 2022). These advancements have significantly accelerated PDD research (Reddy and Zhang 2013).
@@
Based on prior knowledge, molecular optimization methods specifically designed for PDD have not been proposed, primarily due to several challenges in this field. First, encoding molecular interactions with cellular systems differs significantly from existing target-based molecular optimization approaches. Second, the evaluation metrics for these tasks vary, making direct adaptation challenging. While the physicochemical drug properties (e.g., lipophilicity and solubility) can be measured quickly and cost-effectively (Morehead and Cheng 2024; Xiong et al. 2023), and drug-target affinities can be accurately predicted using existing virtual screening techniques (Ru et al. 2022; Lu et al. 2022), these approaches are not directly applicable to PDD. In the context of PDD optimization, the ideal evaluation metric is the IC50 value, which reflects the interaction between molecules and cell lines. Although determining IC50 values experimentally is time-consuming and expensive, they can be predicted by drug response prediction (DRP) models within a certain margin of error. By evaluating the predicted IC50 distribution of an optimized molecular set, we can assess the optimization model’s effectiveness, thereby mitigating the impact of prediction errors.
@@
For this reason, we propose a fragment-masked molecular optimization method for the PDD (FMOP). As shown is Fig. 1, the FMOP method’s optimization conditions include the initial molecule to be optimized, its corresponding masked fragments, the target cell line, and the IC50 value. Notably, the IC50 is an efficacy measure of a single drug response against one specific cell line, obtained through the wet experiment. The output is a batch of optimized molecules under specified conditions. In addition, FMOP employs a regression-free diffusion model to conditionally optimize the molecules’ masked regions, effectively generating new molecules with similar scaffolds and improving IC50 without training. Specifically, molecules’ fragment masks are primarily based on scaffolds and side chains in the molecular structure. First, we apply rule-based constraints to the fragmentation results. Then, we use a pre-trained diffusion model as the generative prior and adjust the unmasked region sampling process during the reverse diffusion iteration using the given drug information, without modifying or conditioning the original diffusion model.
@@
On the large-scale drug response dataset GDSCv2 (Yang et al. 2013), we conduct optimization experiments on all 945 cell lines, demonstrating an in-silico optimization success rate of 94.4% and an average efficacy increase of 5.3%. Additionally, through extensive ablation and visualization experiments, we further demonstrate that FMOP is an effective and robust molecular optimization method with broad application prospects in PDD. This paper’s contributions are as follows:
@@
    Optimization experiments were conducted across all 945 cell lines in the GDSCv2 dataset, demonstrating a 94.4% success rate and a 5.3% average increase in efficacy through optimization. Extensive visualization evaluations further indicate FMOP’s robustness and broad applicability.
Related Work
Molecular optimization aims to improve drug properties, including physicochemical (e.g., solubility, stability, and absorption) and biomedical attributes (e.g., toxicity, target affinity, and drug-cell response), thereby enhancing drug candidate effectiveness.
Existing molecular optimization methods can be broadly categorized into rule- and deep learning-based methods. Rule-based methods, such as pharmacophore modeling, rely on predefined structural rules to suggest modifications. In comparison, deep learning-based methods, such as those utilizing the denoising diffusion probabilistic model (DDPM) (Xie et al. 2024; Peng et al. 2022; Ho, Jain, and Abbeel 2020), have demonstrated great potential for molecular optimization. These methods integrate DDPM’s ability to generate high-quality and diverse molecules with specific optimization goals and conditions to efficiently refine molecular structures and properties.
Despite significant progress, current molecular optimization techniques encounter limitations in meeting PDD requirements. Numerous studies have been proposed optimizing physicochemical properties (Xiong et al. 2023; Born and Manica 2023), such as toxicity and target affinity (Huang et al. 2024; Jorgensen 2009), demonstrating significant potential in accelerating molecular optimization. However, PDD molecular optimization methods have not been thoroughly investigated. Some drugs exhibit favorable solubility and stability in vitro but fail to deliver the expected efficacy in practical applications (Rocha et al. 2023). This gap often occurs due to current methods disregarding the complexity of cellular environments. In addition, current optimization methods are often limited to specific masks or predicting functional group placements (de Souza Neto et al. 2020). These fragment-based approaches often regard molecular optimization as the addition or deletion of individual atoms or chemical bonds (Huang et al. 2024; Cheng et al. 2023). Consequently, when tackling novel mask types or tasks, these methods may struggle to handle complex masking scenarios due to limited training data, thereby restricting their applicability (Bröker et al. 2022).
@@
Method
Problem Formulation. Molecular optimization aims to enhance a molecule’s properties to reveal improved alternatives. For the PDD-based optimization task, the molecule’s property is its efficacy in a specific cell line, denoted as IC50.
Let one molecule be represented as $\mathbf{G}$, the cell line as $C=\left\{c_{1},c_{2},...,c_{m}\right\}$, and their IC50 as $Y=\left\{y_{1},y_{2},...,y_{m}\right\}$. The optimized molecule’s $Y$ with respect to $C$ is denoted as $Y^{\prime}$. Therefore, for a given $(\mathbf{G},c_{i},y_{i})$, the optimized result $y^{\prime}$ should satisfy $y^{\prime}<y$. Based on our fragment-masked method, we decomposed $\mathbf{G}$ into a scaffold $\mathcal{S}_{\mathbf{f}}$ and a side chain $\mathcal{S}_{\mathbf{c}}$, where a mask marks the side chain’s fragment structure. For one molecule $\mathbf{G}=(A,X)$, $\mathcal{M}^{X}\in\mathbb{R}^{\left|X\right|}$ denotes an ordered array of the atom mask $X\in\mathbb{R}^{\left|X\right|}$, where the atoms on the scaffold are labeled as 0 and those on the side chain as 1. In addition, $\mathcal{M}^{A}\in\mathbb{R}^{\left|X\right|\times\left|X\right|}$ represents an ordered array of the edge mask $A\in\mathbb{R}^{\left|X\right|\times\left|X\right|}$, with edges on the side chain labeled as 1 and those on the scaffold or between scaffolds and side chains labeled as 0.
Overview. To generate molecules with a specific distribution under numerical drug response conditions, we employed a regressor-free conditional diffusion method. As illustrated in Fig. 2, we integrated specific conditions about the cell line $C$ and IC50 $Y$ into the scoring estimation to guide the diffusion model. Specifically, to establish the molecules’ conditional constraints, we split the molecule into its scaffold and side chain to generate the fragment masks. Then, we performed molecule splitting according to the Murcko scaffold method.
@@
Molecular Conditional Generation
Typically, the input conditions for molecular optimization methods are categorical. To more precisely and efficiently optimize molecular graphs to specific conditional distributions, our model for conditional generation follows the regressor-free molecular generation method (Li et al. 2024a), which can effectively generate molecules under the given numerical conditions. Moreover, the conditions comprise text labels for the PDD task (i.e., the cell line name and IC50 values) and fragment mask arrays $\mathcal{M}^{X}$ and $\mathcal{M}^{A}$. To effectively receive text conditions, we employed a contrastive learning strategy to align the two feature types. The drug, cell line, and fusion drug response encoders are denoted as $\Phi_{\mathrm{G}}$, $\Phi_{\mathrm{C}}$, and $\Phi_{\mathrm{F}}$ respectively.
Moreover, the text encoder that describes the reaction process between the drugs and cell lines is denoted as $\Phi_{\mathrm{T}}$, and its input text $\Phi_{\mathrm{T}}$ is generated through a template with two parameters ($C,Y$), as the response value of the drug with the $-[]𝐧𝐚𝐦𝐞𝐨𝐟𝐭𝐡𝐞𝐜𝐞𝐥𝐥𝐥𝐢𝐧𝐞\mathbf{\color[rgb]{0.24609375,0.4453125,0.6875}{\left[name\;of\;the\;cell\;line\right]}}$ is $\mathbf{{\color[rgb]{0.24609375,0.4453125,0.6875}{\left[\mathbf{IC}_{50}\right]}}}$. For the $i$-th representations $(d_{i},c_{i})$ generated by the $\Phi_{\mathrm{F}}$ and the $j$-th captions $(c_{j},y_{j})$ produced by the caption encoder in a batch $\mathcal{B}$, we normalized the feature vectors in a hyper-sphere using $u_{i}:=\frac{\Phi_{\mathrm{F}}(d_{i},c_{i})}{\left\|\Phi_{\mathrm{F}}(c_{i})\right\|}$ and $v_{j}:=\frac{\Phi_{\mathrm{T}}(c_{j},y_{j})}{\left\|\Phi_{\mathrm{T}}(c_{j},y_{j})\right\|}$. Finally, the similarity between $u_{i}$ and $v_{j}$ was calculated as $u_{i}^{\mathrm{T}}v_{j}$. Hence, the supervised contrastive loss function $\mathcal{L}_{\mathrm{NCE}}$ can be expressed as:

### v3

# Fragment-Masked Diffusion for Molecular Optimization

@@ Abstract
Molecular optimization is a crucial aspect of drug discovery, aimed at refining molecular structures to enhance drug efficacy and minimize side effects, ultimately accelerating the overall drug development process. Many molecular optimization methods have been proposed, significantly advancing drug discovery. These methods primarily on understanding the specific drug target structures or their hypothesized roles in combating diseases. However, challenges such as a limited number of available targets and a difficulty capturing clear structures hinder innovative drug development. In contrast, phenotypic drug discovery (PDD) does not depend on clear target structures and can identify hits with novel and unbiased polypharmacology signatures. As a result, PDD-based molecular optimization can reduce potential safety risks while optimizing phenotypic activity, thereby increasing the likelihood of clinical success. Therefore, we propose a fragment-masked molecular optimization method based on PDD (FMOP). FMOP employs a regression-free diffusion model to conditionally optimize the molecular masked regions, effectively generating new molecules with similar scaffolds. On the large-scale drug response dataset GDSCv2, we optimize the potential molecules across all 985 cell lines. The overall experiments demonstrate that the in-silico optimization success rate reaches 95.4%, with an average efficacy increase of 7.5%. Additionally, we conduct extensive ablation and visualization experiments, confirming that FMOP is an effective and robust molecular optimization method. The code is available at: https://anonymous.4open.science/r/FMOP-98C2.
Index Terms:
Molecular optimization, fragment-masked, diffusion model, phenotypic drug discovery, drug discovery.
I Introduction
Molecular optimization plays a crucial role in drug discovery [1], which involves the modification and improvement of lead compounds identified through initial screening to enhance their drug-like properties [2]. Historically, molecular optimization is planned manually according to knowledge and experience in the pharmacology, and optimized through fragment-based screening or synthesis [3]. However, manual molecular optimization is not easily scalable to different needs and cannot be automated for large-scale optimization; thus, this strategy is insufficient for meeting the demands of current drug discovery [4].
In recent years, deep learning (DL) methods, particularly diffusion models [5], have been observed to effectively optimize molecules that meet specific conditions [6, 7], with the potential to accelerate traditional paradigms. Molecular optimization can rapidly identify potential drug candidates using existing experimental data and molecular structures, reducing the need for blind experimentation and enhancing research efficiency [8].
@@ I Introduction
Several molecular optimization methods have been proposed to enhance candidate molecule performance, such as target pocket- [9] and property-based molecular optimization [10, 11, 12, 13]. Target pocket-based molecular optimization methods rely on understanding specific pocket structures and their hypothesized roles in combating diseases [14]. Challenges such as a limited number of available targets and difficulty capturing clear structures hinder innovative drug development. In contrast, phenotypic drug discovery (PDD) [15] does not depend on well-defined target structures and can identify hit compounds with novel and multi-target properties. PDD emphasizes the phenotypic effects of molecules within disease-related biological systems [16] and has significantly contributed to the discovery of first-in-class drugs [17]. By observing the phenotypic changes that molecules induce in cells, tissues, or organisms, PDD identifies potential drugs without requiring prior knowledge about specific targets [18]. As a result, PDD’s potential as a drug discovery tool that addressing complex diseases that are not yet fully understood by the scientific community is evident. Extensive PDD research has been conducted, with significant efforts made in constructing relevant datasets such as the genomics of drug sensitivity in cancer (GDSC) [19, 20]. Based on these datasets, numerous artificial intelligence-driven methods [21, 22, 23] have been proposed to predict drug responses against specific cell lines, typically quantified by the half-maximal inhibitory concentration (IC50) [24]. These advancements have significantly accelerated PDD research [25].
@@
Based on prior knowledge, molecular optimization methods specifically designed for PDD have not been proposed, primarily due to several challenges in this field. First, encoding molecular interactions with cellular systems differs significantly from existing target- and properties-based molecular optimization methods. Second, the evaluation metrics for these tasks vary, making direct adaptation challenging. While the physicochemical drug properties (e.g., lipophilicity and solubility) can be measured quickly and cost-effectively [26], and drug-target affinities can be accurately predicted using existing virtual screening techniques [27, 28], these methods are not directly applicable to PDD. In the context of PDD optimization, the ideal evaluation metric is the IC50 value, which reflects the interaction between molecules and cell lines. Although determining IC50 values experimentally is time-consuming and expensive, they can be predicted by drug response prediction (DRP) models within a certain margin of error. By evaluating the predicted IC50 distribution of an optimized molecular set, we can assess the optimization model’s effectiveness, thereby mitigating the impact of prediction errors.
@@
For this reason, we propose a fragment-masked molecular optimization method for the PDD (FMOP). As shown is Fig. 1, the FMOP method’s optimization conditions include the initial molecule to be optimized, its corresponding masked fragments, the target cell line, and the IC50 value. Notably, the IC50 is an efficacy measure of a single drug response against one specific cell line, obtained through the wet experiment. The output is a batch of optimized molecules under specified conditions. FMOP employs a regression-free diffusion model to conditionally optimize the molecules’ masked regions, effectively generating new molecules with similar scaffolds and improving IC50. Specifically, molecules’ fragment masks are primarily based on scaffolds and side chains in the molecular structure. First, we apply rule-based constraints to the fragmentation results. Then, we use a pre-trained diffusion model as the generative prior and adjust the unmasked region sampling process during the reverse diffusion iteration using the given drug information, without modifying or conditioning the original diffusion model.
@@
On the large-scale drug response dataset GDSCv2 [19], we conducted optimization experiments on all 985 cell lines, demonstrating an in-silico optimization success rate of 95.4% and an average efficacy increase of 7.5%. It is important to emphasize that the optimization task across 985 cell lines is analogous to optimization across different properties. Our method requires training only once to cover all tasks, whereas other methods would need to train independently for each of the 985 cell lines. Additionally, through extensive ablation and visualization experiments, we further demonstrate that FMOP is an effective and robust molecular optimization method with broad application prospects in PDD. This paper’s contributions are as follows:
@@
    Optimization experiments were conducted across all 985 cell lines in the GDSCv2 dataset, demonstrating a 95.4% success rate and a 7.5% average increase in efficacy through optimization. Extensive visualization evaluations further indicate FMOP’s robustness and broad applicability.
II Related Work
Molecular optimization aims to improve drug properties, including physicochemical (e.g., solubility, stability, and absorption) and biomedical attributes (e.g., toxicity, target affinity, and drug-cell response) [29, 30, 31], thereby enhancing drug candidate effectiveness. Existing molecular optimization methods could be broadly categorized into rule- and deep learning-based methods. Rule-based methods [32, 13], such as pharmacophore modeling and fragment libraries built using JT-VAE [33], relied on predefined structural rules to suggest modifications. In comparison, deep learning-based methods [34, 35], such as those utilizing the denoising diffusion probabilistic model (DDPM) [36, 37, 5], demonstrated great potential for molecular optimization. More recent methods were flow-based [38], graph-based [32, 39, 40, 41, 12] and Transformer-based [10], integrating diffusion models to enhance optimization outcomes [42]. These methods integrated DDPM’s ability to generate high-quality and diverse molecules with specific optimization goals and conditions to efficiently refine molecular structures and properties.
Despite significant progress, current molecular optimization techniques encountered limitations in meeting PDD requirements. Numerous studies have been proposed optimizing physicochemical properties [43, 44, 45], such as toxicity and target affinity [9, 46], demonstrating significant potential in accelerating molecular optimization. However, PDD molecular optimization methods have not been thoroughly investigated. Some drugs showed favorable solubility and stability in vitro but failed to deliver the expected efficacy in practical applications [47]. This gap occurred due to current methods disregarding the complexity of cellular environments. In addition, current optimization methods were often limited to specific masks or predicting functional group placements [48, 12, 32]. These fragment-based methods regarded molecular optimization as the addition or deletion of individual atoms or chemical bonds [9, 49, 50]. Consequently, when tackling novel mask types or tasks, these methods may struggle to handle complex masking scenarios due to limited training data, thereby restricting their applicability [51]. Most importantly, the training and optimization mechanisms employed by these methods were originally designed to train only on a set of molecules with a specific property. When the properties change, these models have to be retrained, so the requirements of modeling multiple properties at the same time and optimizing the molecule for different properties could not be met if secondary training was required.
@@ II Related Work
III Method
Problem Formulation. Molecular optimization aims to enhance a molecule’s properties to reveal improved alternatives. For the PDD-based optimization task, the molecule’s property is its efficacy in a specific cell line, denoted as IC50. Table I provides a summary of the notation used in this paper.
Let one molecule be represented as $\mathbf{G}$, the cell line as $\mathbf{C}=\left\{c_{1},c_{2},...,c_{m}\right\}$, and their IC50 as $\mathbf{Y}=\left\{y_{1},y_{2},...,y_{m}\right\}$. The optimized molecule’s $\mathbf{Y}$ with respect to $\mathbf{C}$ is denoted as $\mathbf{Y}^{\prime}$. Therefore, for a given $(\mathbf{G},c_{i},y_{i})$, the optimized result $y^{\prime}$ should satisfy $y^{\prime}<y$. Based on our fragment-masked method, we decomposed $\mathbf{G}$ into a scaffold $\mathcal{S}_{\mathbf{f}}$ and a side chain $\mathcal{S}_{\mathbf{c}}$, where a mask marks the side chain’s fragment structure. For one molecule $\mathbf{G}=(A,X)$, $\mathcal{M}^{X}\in\mathbb{R}^{\left|X\right|}$ denotes an ordered mask matrix of the atom matrix $X\in\mathbb{R}^{\left|X\right|}$, where the atoms on the scaffold are labeled as 0 and those on the side chain as 1. In addition, $\mathcal{M}^{A}\in\mathbb{R}^{\left|X\right|\times\left|X\right|}$ represents an ordered mask matrix of the edge matrix $A\in\mathbb{R}^{\left|X\right|\times\left|X\right|}$, with edges on the side chain labeled as 1 and those on the scaffold or between scaffolds and side chains labeled as 0.
TABLE I: Mathematical Notations
Notations Descriptions
$\mathbf{G}$ Molecule graph
$A$ Edge adjacency matrix
$X$ Node feature matrix
$\mathbf{C}=\{c_{1},c_{2},...,c_{m}\}$ Set of cell lines
$\mathbf{Y}=\{y_{1},y_{2},...,y_{m}\}$ IC50 values
$\mathbf{Y}^{\prime}=\{y_{1}^{{}^{\prime}},y_{2}^{{}^{\prime}},...,y_{m}^{{}^{\prime}}\}$ Optimized IC50 values
$\mathcal{S}_{\mathbf{f}},\mathcal{S}_{\mathbf{c}}$ The scaffold and side chain from one molecule
$\mathbf{c}=(c_{i},y_{j})$ Sampling condition
$\mathcal{M}^{A},\mathcal{M}^{X}$ Mask matrix (edge and node)
$A^{\text{ukn}},X^{\text{ukn}}$ Unknown regions mask matrix (edge and node)
$A^{\text{kn}},X^{\text{kn}}$ Known regions mask matrix (edge and node)
Overview. To generate molecules with a specific distribution under numerical drug response conditions, we employed a regressor-free conditional diffusion method. As illustrated in Fig. 2, we integrated specific conditions about the cell line $\mathbf{C}$ and IC50 $\mathbf{Y}$ into the scoring estimation to guide the diffusion model. Specifically, to establish the molecules’ conditional constraints, we split the molecule into its scaffold and side chain to generate the fragment masks. Then, we performed molecule splitting according to the Murcko scaffold method.
@@ III Method
III-A Molecular Conditional Generation
Typically, the input conditions for molecular optimization methods are categorical. To more precisely and efficiently optimize molecular graphs to specific conditional distributions, our model for conditional generation follows the regressor-free molecular generation method [52], which can effectively generate molecules under the given numerical conditions. Only the cell line type needs to be adjusted when the optimization target changes. Various attribute conditions, such as cell line types, have been fully incorporated during the training phase, and different attributes are unified through contrastive learning using a contrastive learning strategy. Moreover, the conditions comprise text labels for the PDD task (i.e., the cell line name and IC50 values) and fragment mask arrays $\mathcal{M}^{X}$ and $\mathcal{M}^{A}$. To effectively receive text conditions, we employed a contrastive learning strategy to align the two feature types. The drug, cell line, and fusion drug response encoders are denoted as $\Phi_{\mathrm{G}}$, $\Phi_{\mathrm{C}}$, and $\Phi_{\mathrm{F}}$ respectively.
Moreover, the text encoder that describes the reaction process between the drugs and cell lines is denoted as $\Phi_{\mathrm{T}}$, and its input text $\Phi_{\mathrm{T}}$ is generated through a template with two parameters ($\mathbf{C},\mathbf{Y}$), as the response value of the drug with the $-[]𝐧𝐚𝐦𝐞𝐨𝐟𝐭𝐡𝐞𝐜𝐞𝐥𝐥𝐥𝐢𝐧𝐞\mathbf{\color[rgb]{0.24609375,0.4453125,0.6875}{\left[name\;of\;the\;cell\;line\right]}}$ is $\mathbf{{\color[rgb]{0.24609375,0.4453125,0.6875}{\left[\mathbf{IC}_{50}\right]}}}$. For the $i$-th representations $(d_{i},c_{i})$ generated by the $\Phi_{\mathrm{F}}$ and the $j$-th captions $(c_{j},y_{j})$ produced by the caption encoder in a batch $\mathcal{B}$, we normalized the feature vectors in a hyper-sphere using $u_{i}:=\frac{\Phi_{\mathrm{F}}(d_{i},c_{i})}{\left\|\Phi_{\mathrm{F}}(c_{i})\right\|}$ and $v_{j}:=\frac{\Phi_{\mathrm{T}}(c_{j},y_{j})}{\left\|\Phi_{\mathrm{T}}(c_{j},y_{j})\right\|}$. Finally, the similarity between $u_{i}$ and $v_{j}$ was calculated as $u_{i}^{\mathrm{T}}v_{j}$. Hence, the supervised contrastive loss function $\mathcal{L}_{\mathrm{NCE}}$ can be expressed as: